5. Solve. 3x + 1 + 1 = x
Answer #1

ل = (+دان | + ر+د3 Ans ۱+دن 3 = H - cuben both side cube 3 هن فرد ہلی = ۸۱ مقاله 3 هدى - ز - 3 - ( ره) را در . من Am 6 = 2- د